BOX I CONTAINS 5 RED AND 6 BLUE BALLS. BOXII CONTAINS 6 RED AND 4 BLUE BALLS, A BALL IS DRAWN AT RANDOM FROM BOX I AND TRANSFERRED TO BOX II. NOW FROM BOX II A BALL IS DRAWN AT RANDOM. WHAT IS THE PROBABLITY THAT IS RED?
Answer by jorel1380(3719)

The probability that the first ball drawn was red is 5/11. The probability that the next ball drawn is red is (5/11+6)/11=(71/11)/11=71/121. ☺☺☺☺